OpenStudy (anonymous):

The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day?

OpenStudy (welshfella):

you should have the confidence to so these on your own now

OpenStudy (anonymous):

you got this from flvs

OpenStudy (anonymous):

so before we start we have to know what type of equation this is

OpenStudy (anonymous):

the type of brackets are a dead give away because it shows us it is a absolute value equation

OpenStudy (anonymous):

since we now know it is a absolute value equation we know you can only have 2 answers

OpenStudy (anonymous):

When dealing with absolute equations you know you are going to have a + and a -. For instance 3|x − 8| + 15 = 18 Subtract 15 from both sides of the equal sign. 3|x − 8| + 15 = 18 3|x − 8| + 15-15 = 18 - 15 3|x − 8| = 3 Now divide 3|x − 8| = 3 by three on each side of the equal sign. 3|x − 8| /3 = 3 /3 |x − 8| = 1 Now we have |x − 8| = 1 but we also have |x − 8| = -1 notice the - 1. When dealing with absolute equations like this you will have a + and a - Now just solve these two equations |x − 8| = 1 and |x − 8| = -1

OpenStudy (anonymous):

why did you give the answer

OpenStudy (anonymous):

@Matt_J02 I did not. He / she still needs to finish it.

OpenStudy (anonymous):

I helped show her or him how to break it down because he or she might be having trouble in breaking down the equation to a form she or he can work with.

OpenStudy (welshfella):

3|x − 8| + 15 = 18 subtract 15 from both sides 3|x - 8| = 3 divide both sides by 3 |x - 8| = 1 solve the 2 equations x - 8 = 1 and x - 8 = -1

OpenStudy (welshfella):

@Nixy - you need to remove the mod lines around the x - 8

OpenStudy (anonymous):

I was leaving that for them to do :-) @welshfella

OpenStudy (welshfella):

nevertheless it could confuse them at that stage

OpenStudy (anonymous):

@welshfella maybe but I would hope they can at least remove them at this stage in their study.

OpenStudy (welshfella):

- oh ok - i guess it was a deliberate mistake on your part.

OpenStudy (anonymous):

@welshfella Not a mistake at all. Lets just take it a step further so he or she is not perplexed. I mean if he or she cannot remove absolute bars I am sure he or she cannot solve x-8 = 1 and x-8 = -1 x-8 = 1 and x-8 = -1 Now add 8 to each side of the equation |x-8| = 1 x-8 = 1 x - 8+8 =1 + 8 x = 9 |x-8| = -1 x-8 = -1 x-8+8 = -1+8 x= 7

OpenStudy (welshfella):

nevertheless writing |x - 8! = -1 is an incorrect statement

OpenStudy (anonymous):

@welshfella it is not if you remove the bars, which is the simplest thing to do in this problem....I was showing the user a process . Nitpick much? Also, you stated |x - 8! = -1. You should use a bar | and not an ! mark. I might not understand what you mean when you stat |x - 8! = -1 ...... Rolls eyes

OpenStudy (welshfella):

its not nitpicking at all and typos are easily made the fact is x - 8 = -1 does not follow from |x - 8| = -1

OpenStudy (anonymous):

@welshfella ok, whatever. Have a good day........Roll eyes